Please show me how to solve this system of equations using the substitution method.
3x-11y=2
x-5y=2
Thanks

Answer by Alan3354(30983)

Please show me how to solve this system of equations using the substitution method.
3x-11y=2
x-5y=2 --> x = 5y+2
Sub into the 1st eqn
3x-11y=2
3*(5y+2) - 11y = 2
15y+6 - 11y = 2
4y+6 = 2
4y = -4
y = -1
-----------
x = 5y+2
x = -3